How many three-digit numbers can be formed using elements from the set if no element may be used more than once in a number and the number must be odd?

Found 2 solutions by stanbon, sudhanshu_kmr:
Answer by stanbon(75887)

How many three-digit numbers can be formed using elements from the set if no element may be used more than once in a number and the number must be odd?
---------------
Hundreds digit: 9 ways
Tens digit: 9 ways
Ones digit: 5 ways
-----------------------
Total number: 9*9*5 = 405
===============================
Cheers,
Stan H.
===========

Answer by sudhanshu_kmr(1152)


no. of ways to put an odd digit at unit place = 5 [i.e 1,3,5,7 or 9]

no. of ways to put any other digit at 100's place = 8
( except '0' and that digit which at unit place)

similarly, no. of ways to put any digit at 10's place = 8
( any digit except at unit place and 100's place)


total no. of ways = 5*8*8 = 320